OpenStudy (anonymous):

(x+6)^2 Is this right? (x+6)^2 =36x

OpenStudy (lynfran):

im guessing you have to expand this...??

OpenStudy (anonymous):

Not really sure if i do or not all they gave me was (x+6)^2?

OpenStudy (solomonzelman):

Have you ever done problems that asked you "to expand" an expression?

OpenStudy (anonymous):

Nope.

OpenStudy (solomonzelman):

Lets start from this. \(\large\color{black}{\rm \displaystyle expression\text{#}1(expression\text{#}2) }\) is same thing as \(\large \color{black}{\rm expression\text{#}1~\cdot~expression\text{#}2}\) you know that right?

OpenStudy (anonymous):

? What is expression 1 and 2 suppose to mean?

OpenStudy (solomonzelman):

any expreesions, i am just naming them like this.

OpenStudy (solomonzelman):

this is equivalent to saying a(b) is same as saying a•b

OpenStudy (anonymous):

Okay.

OpenStudy (solomonzelman):

good.

OpenStudy (solomonzelman):

So, if I say a(b), then I know that it is a•b and it would be the same thing as a•b.

OpenStudy (solomonzelman):

(Just trying to get general way of writing things straight)

OpenStudy (solomonzelman):

a(b) = a•b = ab

OpenStudy (solomonzelman):

all 3 parts are same...

OpenStudy (anonymous):

Alright

OpenStudy (solomonzelman):

now, if I said: a(b+c), then what am I really doing? a(b+c) = a•(b+c) = a•b+a•c=ab+ac

OpenStudy (solomonzelman):

I am multiplying everything inside the parenthesis times a.

OpenStudy (solomonzelman):

is this making some sense?

OpenStudy (anonymous):

Yeah, just not sure what this has to do with the problem i asked.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

Anyone?

OpenStudy (jdoe0001):

\(\large (x+6)^2 \implies (x+6)(x+6)\implies \begin{cases} x\cdot x\implies &?\\ x\cdot 6\implies &?\\ 6\cdot x\implies &?\\ 6\cdot 6\implies &? \end{cases}\)

OpenStudy (anonymous):

Okay think i got it. x*x = X^2, x*6 = x^6, 6*x= 6x and 6*6 = 36

jimthompson5910 (jim_thompson5910):

(x+6)^2 is the same as (x+6)*(x+6) squaring something means you're multiplying it by itself To expand out (x+6)*(x+6) you can use the FOIL rule but I like a table better because you can use a table for things that FOIL can't do